Financial analysts have estimated the returns on shares of Drucker Corporation and the overall market portfolio under various economic conditions as follows. The return for Drucker in the following three economic states of nature are forecasted to be:-18% in recession, +9% in moderate growth, and +36% in a boom. Estimates for the market as a whole in the same economic states are-12% in recession, +6% in moderate growth, and +22% in boom. The analyst considers each state to be equally likely. Using these data, compute the beta of Drucker Corporations stock. Place your answer as a number with at least three decimal places
Answer #1
ANALYSIS OF BETA
Market condition Probability (P) Return of stock (Rs) Return of Market(Rm) Ds=Rs-avg.Rs Dm=Rm-avg.Rm Ds^2 Dm^2 P*Dm^2 P*Ds^2
Recession 0.333 -18 -12 -27 -17.33 729 300.4444 100.1481 243
Moderate Growth 0.333 9 6 0 0.67 0 0.444444 0.148148 0
Boom 0.333 36 22 27 16.67 729 277.7778 92.59259 243
Total 1.000 1458 578.6667 192.8889 486
Avg.Rs 9
Avg.Rm 5.33
Variance of Market P*Dm^2 192.89
Std Deviation (Rs) 22.04540769
Std Deviation (Rm) 13.88844444
Coorelation (Cor) 0.9994 Approx.=+1
Covariance Cor*Std.dev(Rs)*Std Dev(Rm)
Beta of Security Covarianc(stock, market)/Variance of Market `=(22.0454*13.8844*1)/192.89
Answer 1.587